# OAT-FM: Optimal Acceleration Transport for Improved Flow Matching

### v1

@@ C.2 Low-Dimensional OT Benchmark
Dataset. Following the prior work in (Tong et al., 2024), we evaluate performance on five two-dimensional distribution mapping tasks. These benchmarks test the ability of a model to learn a transport map from a source distribution to the target distribution. The specific pairs are: $i)$ a standard Gaussian to a mixture of 8 Gaussians ($\mathcal{N}\to 8$gs), ii)ii) a standard Gaussian to two interleaved moons ($\mathcal{N}\to$ moons), $iii)$ a standard Gaussian to an S-shaped curve ($\mathcal{N}\to$ scurve), $iv)$ moons to 8 Gaussians (moons $\to 8$gs), and $v)$ 8 Gaussians to moons (88gs $\to$ moons).
@@
Training. For all experiments, the vector field $v_{\theta}$ is parameterized by a standard Multi-Layer Perceptron (MLP) that accepts concatenated position and time vectors as input. The MLP consists of 3 hidden layers, each with a width of 64 neurons, followed by a SELU activation function. The final layer maps the representation back to a 2-dimensional vector, representing the velocity at the given point in space-time. First, we train the baseline models, i.e., FM (Lipman et al., 2023), I/SB/OT-CFM (Tong et al., 2024), and VP-CFM Albergo et al. (2023), for 20,000 batches (each batch contains 256 data points). Subsequently, each of these pre-trained models is refined using our OAT-FM for an additional 20,000 batches. During the OAT-FM training, we employ a slowly-updating strategy. The OAT-FM is updated with a hard copy of the online model’s weights every 500 batches.
@@
Evaluation. We assess model performance using two key metrics. First, to measure the quality of the learned terminal distribution, we compute the 2-Wasserstein distance ($\mathcal{W}_{2}$) between the generated samples and the true target samples. Second, to evaluate the efficiency of the learned transport path, we use the Normalized Path Energy (NPE), defined as $\text{NPE}=\frac{|\text{PE}(v_{\theta})-\mathcal{W}_{2}^{2}(\rho_{0},\rho_{1})|}{\mathcal{W}_{2}^{2}(\rho_{0},\rho_{1})}$, where $\text{PE}(v_{\theta})=\mathbb{E}_{x_{0}\sim\rho_{0}}[\int_{0}^{1}\|v_{\theta}(t,x_{t})\|^{2}\mathrm{d}t]$ is the path’s kinetic energy and $\mathcal{W}_{2}^{2}(\rho_{0},\rho_{1})$ is the ground-truth squared Wasserstein distance between the source and target distributions. An NPE value near zero indicates that the learned path is close to the dynamic optimal transport plan. For both metrics, we use a test set of 1,024 samples. Trajectories are generated by integrating the learned vector field from $t=0$ to $t=1$ using a 4th-order Runge-Kutta (RK4) solver (Dormand and Prince, 1980) with 101 discretization steps and absolute and relative error tolerances of $10^{-6}$. The path energy integral is numerically approximated using the trapezoidal rule over the computed trajectory points.
@@ C.3 CIFAR-10 $\times32$
Dataset. We evaluate our model on the CIFAR-10 dataset (Krizhevsky, 2009), a widely-used benchmark for image generation. The dataset consists of 60,000 color images with size $\times32$ in 10 classes, partitioned into 50,000 training images and 10,000 test images.
Training. We refine open-source FM and I/OT-CFM models from https://github.com/atong01/conditional-flow-matching/tree/main/examples/images/cifar10 that were trained for 400K iterations. The neural network is a U-Net architecture Ho et al. (2020) with an implementation from the guided-diffusion repository in https://github.com/openai/guided-diffusion. The U-Net has a base of 128 channels, channel multipliers of $[1,2,2,2]$, two residual blocks per resolution, and applies four-head self-attention at the $\times16$ resolution. The boundary velocity fields required for the OAT-FM loss are estimated using a target network, which is an exponential moving average (EMA) of the online model’s weights with a decay of 0.9999 (See Algorithm 1). We , we use the Adam optimizer with a learning rate of $2\times 10^{-4}$, and apply gradient clipping with a maximum norm of 1.0. We use OAT-FM to train these three models with an additional 1K iterations. For EDM, we downloaded the checkpoints from https://drive.google.com/drive/folders/18dWE-LiodXdCG0RDNegySzRnyRdcwamW. This model utilizes the DDPM++ architecture (SongUNet) from the work in (Song et al., 2021a), which is a U-Net composed of residual blocks, self-attention, and positional timestep embeddings. Following the methodology in (Lee et al., 2024) to convert a pre-trained score model into a flow model (as described in Appendix C.1), the network is wrapped in a velocity-prediction head. This wrapper adapts the model’s output to predict a velocity field, making it directly compatible with our flow matching objective. Using OAT-FM, it is trained for an additional 12K iterations in EMA strategy, with a learning rate warmup over the first 1K iterations from 3e-5 to 3e-4.
Evaluation. In the inference phase, We employ the adaptive Dopri5 solver for FM, I-CFM, and OT-CFM, and the Heun solver with 35 steps for EDM. For each model, we report image quality measured by the Fréchet Inception Distance (FID) (Heusel et al., 2017), along with the number of training iterations (#Iter.) and the number of function evaluations (NFE) during inference. In addition, we also present the results of FM, I-CFM, and OT-CFM using the Euler solver with 100 steps in Table 7.
@@ C.4 ImageNet $\times256$
Dataset. We extend our evaluation to class-conditional generation on the ImageNet $\times256$ benchmark (Deng et al., 2009). This dataset, a standard for large-scale image generation, consists of approximately 1.28 million training images, categorized into 1,000 classes.
Training. Our generative model for ImageNet is a Scalable Interpolant Transformer (SiT) (Ma et al., 2024), which utilizes the Diffusion Transformer (DiT) (Peebles and Xie, 2023) backbone. We downloaded the checkpoint from https://github.com/willisma/SiT. The model operates in the latent space of a pre-trained variational autoencoder (VAE) (Kingma and Welling, 2013). The network architecture is the XL/2 version of SiT. The interpolant framework allows for flexible choices of path-type and model prediction targets, with linear path and velocity prediction being the default configuration. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and no weight decay. The target network is updated via an EMA strategy of the online model’s weights with a decay of 0.9999. When applying Algorithm 1, to prevent cross-class interference within mini-batches, our sampling strategy assigns a unique class to each GPU for every training iteration. This ensures that each local batch consists solely of images from a single class. To incorporate Classifier-Free Guidance (CFG), we follow the training protocol of SiT-XL and randomly drop class labels with a probability of 0.1. This procedure partitions each mini-batch into two subsets: a conditional (labeled) group and an unconditional (unlabeled) group. Our method computes the OAT plan and performs the corresponding sample pairing independently within each subset. Finally, the paired samples from both groups are combined to compute the training loss. The pre-trained SiT-XL is trained by OAT-FM with additional 48K iterations or 5 epochs.
Evaluations. We employ both ODE and SDE samplers for evaluations. For ODE-based sampling, we utilize the adaptive step-size Dopri5 solver (Dormand and Prince, 1980), configured with an absolute tolerance of $1\times 10^{-6}$ and a relative tolerance of $1\times 10^{-3}$. For SDE-based sampling, we use a fixed-step Euler-Maruyama solver with 250 steps. In both settings, we leverage classifier-free guidance to improve sample quality. The results achieved under different CFG scales are shown in Table 8.
@@
Appendix D More Visual Results
We generate images of different classes by the original SiT-XL and that refined by OAT-FM, respectively. For each example, we start from the same noise point to generate images by the two models. Following existing methods (Ma et al., 2024; Peebles and Xie, 2023), we set the CFG scale to be 4.0 for good visual effects. Some typical results are shown in Figures 4 and 7, demonstrating that the refinement achieved by OAT-FM indeed leads to better image quality. In some cases, applying OAT-FM just leads to improved image details and maintains the main semantic and spatial content created by SiT-XL. However, in the other cases, applying OAT-FM makes the model creates totally different images compared to those created by SiT-XL. This phenomenon indicates that the coupling associated with OAT-FM is different from that of the original model, which may change the flow trajectories a lot.
@@ Appendix D More Visual Results
(a) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(b) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(c) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(d) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(e) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(f) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(g) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(h) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
Figure 7: Some generation results achieved by the two methods.

### v2

@@ C.2 Low-Dimensional OT Benchmark
Dataset. Following the prior work in (tong2024improving), we evaluate performance on five two-dimensional distribution mapping tasks. These benchmarks test the ability of a model to learn a transport map from a source distribution to the target distribution. The specific pairs are: $i)$ a standard Gaussian to a mixture of 8 Gaussians ($\mathcal{N}\to 8$gs), ii)ii) a standard Gaussian to two interleaved moons ($\mathcal{N}\to$ moons), $iii)$ a standard Gaussian to an S-shaped curve ($\mathcal{N}\to$ scurve), $iv)$ moons to 8 Gaussians (moons $\to 8$gs), and $v)$ 8 Gaussians to moons (88gs $\to$ moons).
@@
Training. For all experiments, the vector field $v_{\theta}$ is parameterized by a standard Multi-Layer Perceptron (MLP) that accepts concatenated position and time vectors as input. The MLP consists of 3 hidden layers, each with a width of 64 neurons, followed by a SELU activation function. The final layer maps the representation back to a 2-dimensional vector, representing the velocity at the given point in space-time. First, we train the baseline models, i.e., FM (lipman2023flow), I/SB/OT-CFM (tong2024improving), and VP-CFM (albergo2023stochastic), for 20,000 batches (each batch contains 256 data points). Subsequently, each of these pre-trained models is refined using our OAT-FM for an additional 20,000 batches. During the OAT-FM training, we employ a slowly-updating strategy. The OAT-FM is updated with a hard copy of the online model’s weights every 500 batches. For the OAT-FM objective, we set the balancing hyperparameter $\alpha$ to 0.70.
@@
Evaluation. We assess model performance using two key metrics. First, to measure the quality of the learned terminal distribution, we compute the 2-Wasserstein distance ($\mathcal{W}_{2}$) between the generated samples and the true target samples. Second, to evaluate the efficiency of the learned transport path, we use the Normalized Path Energy (NPE) defined in (14). An NPE value near zero indicates that the learned path is close to the dynamic optimal transport plan. For both metrics, we use a test set of 1,024 samples. Trajectories are generated by integrating the learned vector field from $t=0$ to $t=1$ using a 4th-order Runge-Kutta (RK4) solver (dormand1980family) with 101 discretization steps and absolute and relative error tolerances of $10^{-6}$. The path energy integral is numerically approximated using the trapezoidal rule over the computed trajectory points.
Comparisons in one-step generation. Table 7 presents our experiments in the one-step generation setting. We include OFM (kornilov2024optimal) as an additional baseline. Here, we only compare different methods on data fitting (2-Wasserstein) because all the methods apply one-step generation. It is worth noting that OFM requires an ICNN architecture, whereas our method and the other baselines utilize the same MLP backbone. The results demonstrate the effectiveness of our method.
Table 7: A comparative analysis of one-step generation performance, evaluating data fitting (2-Wasserstein) and optimal transport approximation (normalized path energy). We run each task in five trials and record the average performance and standard deviation.
[…]
Method $\mathcal{W}_{2}^{2}\downarrow$ $\mathcal{W}_{2}^{2}\downarrow$ $\mathcal{W}_{2}^{2}\downarrow$ $\mathcal{W}_{2}^{2}\downarrow$ $\mathcal{W}_{2}^{2}\downarrow$
OFM $\text{0.71}_{\pm\text{0.27}}$ $\text{0.22}_{\pm\text{0.02}}$ $\text{0.22}_{\pm\text{0.01}}$ $\text{1.99}_{\pm\text{0.25}}$ $\text{0.46}_{\pm\text{0.11}}$
FM $\text{21.50}_{\pm\text{0.07}}$ $\text{4.05}_{\pm\text{0.01}}$ $\text{8.08}_{\pm\text{0.05}}$ $\text{79.01}_{\pm\text{1.18}}$ $\text{13.48}_{\pm\text{0.15}}$
+OAT-FM $\textbf{0.43}_{\pm\text{0.08}}$ $\textbf{0.14}_{\pm\text{0.02}}$ $\textbf{0.12}_{\pm\text{0.02}}$ $\textbf{1.41}_{\pm\text{0.23}}$ $\textbf{0.32}_{\pm\text{0.07}}$
I-CFM $\text{23.46}_{\pm\text{0.15}}$ $\text{4.47}_{\pm\text{0.08}}$ $\text{7.86}_{\pm\text{0.03}}$ $\text{79.31}_{\pm\text{1.41}}$ $\text{14.55}_{\pm\text{0.28}}$
+OAT-FM $\textbf{0.42}_{\pm\text{0.10}}$ $\textbf{0.17}_{\pm\text{0.02}}$ $\textbf{0.12}_{\pm\text{0.01}}$ $\textbf{1.41}_{\pm\text{0.26}}$ $\textbf{0.51}_{\pm\text{0.13}}$
VP-CFM $\text{12.17}_{\pm\text{0.29}}$ $\text{8.46}_{\pm\text{0.25}}$ $\text{3.70}_{\pm\text{0.05}}$ $\text{47.73}_{\pm\text{1.94}}$ $\text{7.51}_{\pm\text{0.21}}$
+OAT-FM $\textbf{0.45}_{\pm\text{0.11}}$ $\textbf{0.20}_{\pm\text{0.03}}$ $\textbf{0.12}_{\pm\text{0.01}}$ $\textbf{1.35}_{\pm\text{0.23}}$ $\textbf{0.36}_{\pm\text{0.07}}$
SB-CFM $\text{0.63}_{\pm\text{0.09}}$ $\text{0.19}_{\pm\text{0.04}}$ $\text{0.20}_{\pm\text{0.04}}$ $\text{1.91}_{\pm\text{0.35}}$ $\text{0.35}_{\pm\text{0.05}}$
+OAT-FM $\textbf{0.42}_{\pm\text{0.10}}$ $\textbf{0.13}_{\pm\text{0.01}}$ $\textbf{0.10}_{\pm\text{0.01}}$ $\textbf{1.22}_{\pm\text{0.26}}$ $\textbf{0.33}_{\pm\text{0.07}}$
OT-CFM $\text{0.48}_{\pm\text{0.07}}$ $\text{0.19}_{\pm\text{0.05}}$ $\textbf{0.12}_{\pm\text{0.01}}$ $\text{1.64}_{\pm\text{0.22}}$ $\text{0.35}_{\pm\text{0.08}}$
+OAT-FM $\textbf{0.40}_{\pm\text{0.08}}$ $\textbf{0.13}_{\pm\text{0.00}}$ $\textbf{0.12}_{\pm\text{0.03}}$ $\textbf{1.23}_{\pm\text{0.19}}$ $\textbf{0.30}_{\pm\text{0.06}}$
@@ C.3 CIFAR-10 $\times32$
C.3.1 Basic Settings
Dataset. We evaluate our model on the CIFAR-10 dataset (krizhevsky2009learning), a widely-used benchmark for image generation. The dataset consists of 60,000 color images with size $\times32$ in 10 classes, partitioned into 50,000 training images and 10,000 test images.
Training. We refine publicly available FM, I-CFM, and OT-CFM models from https://github.com/atong01/conditional-flow-matching/tree/main/examples/images/cifar10 that were pre-trained for 400K iterations. The neural network is a U-Net architecture (Ho2020Denoising) with an implementation adapted from the guided-diffusion repository in https://github.com/openai/guided-diffusion. The U-Net has a base of 128 channels, channel multipliers of $[1,2,2,2]$, two residual blocks per resolution, and applies four-head self-attention at the $\times16$ resolution. The boundary velocity fields required for the OAT-FM loss are estimated using a target network, which is an exponential moving average (EMA) of the online model’s weights with a decay of 0.9999 (See Algorithm 1). For the OAT-FM objective, we set the balancing hyperparameter $\alpha$ to 0.75. We use the Adam optimizer with a learning rate of $2\times 10^{-4}$, and apply gradient clipping with a maximum norm of 1.0. We use OAT-FM to train these three models with an additional 1K iterations. For EDM, we downloaded the checkpoints from https://drive.google.com/drive/folders/18dWE-LiodXdCG0RDNegySzRnyRdcwamW. This model utilizes the DDPM++ architecture (SongUNet) from the work in (Song2021ScoreBased), which is a U-Net composed of residual blocks, self-attention, and positional timestep embeddings. Following the methodology in (lee2024improving) to convert a pre-trained score model into a flow model (as described in Appendix C.1), the network is wrapped in a velocity-prediction head. This wrapper adapts the model’s output to predict a velocity field, making it directly compatible with our flow matching objective. Using OAT-FM, it is trained for an additional 12K iterations in EMA strategy, with a learning rate warmup over the first 1K iterations from 3e-5 to 3e-4.
Evaluation. In the inference phase, We employ the adaptive Dopri5 solver for FM, I-CFM, and OT-CFM, and the Heun solver with 35 steps for EDM. For each model, we report image quality measured by the Fréchet Inception Distance (FID) (heusel2017gans), along with the number of training iterations (#Iter.) and the number of function evaluations (NFE) during inference.
C.3.2 More Experiment Results
Additional Results wth Euler Solvers. In addition, we also present the results of FM, I-CFM, and OT-CFM using the Euler solver with 100 steps in Table 8.
@@ C.3.2 More Experiment Results
Figure 5: Stability analysis of OAT-FM fine-tuning on CIFAR-10. Starting from a pre-trained FM model (400K batches), we refine the model using OAT-FM for an additional 1K training samples, recording metrics every 0.1K samples. We track generation quality (FID with 10 and 100 Euler steps) and flow straightness.
Stability Analysis. To investigate the stability of our method and verify that the OAT-FM refinement does not lead to deterioration of the transport map or distribution drift over time, we conducted a fine-grained quantitative analysis on the CIFAR-10 dataset. We initialize our training from the pre-trained FM model (lipman2023flow) and fine-tune it with OAT-FM on an additional 1K training samples, recording performance metrics per 0.1K samples. As shown in Figure 5, we track three key metrics: generation quality using an Euler solver with 10 steps (FID-10) and 100 steps (FID-100), and the flow straightness score (liu2022rectified; lee2024improving). [^5] The straightness score is computed with 100 integration steps by measuring the mean squared error between the actual velocities along the ODE trajectory and the constant velocity from the initial to the final state, with lower values indicating straighter flows. We observe a consistent, monotonic improvement across all three metrics as the number of OAT-FM training samples increases. This validates the stability of the proposed two-phase training paradigm.
Sensitivity to Phase 1 pre-trained model. As presented in Table 9, we evaluated the efficacy of OAT-FM when applied to pre-trained models at different stages of convergence. Specifically, we utilized checkpoints from FM, I-CFM, and OT-CFM trained for 100K, 200K, 300K, and 400K batches on CIFAR-10. OAT-FM (applied for only 1K batches) consistently improves the generation quality (FID) and reduces the Number of Function Evaluations (NFE) across all initialization points. Notably, even for less-converged models (e.g., FM at 100K batches), OAT-FM successfully reduces the FID from 6.11 to 5.60, demonstrating that our method does not require a near-perfect velocity field to yield benefits. This indicates that OAT-FM is robust to the quality of the initial velocity estimates and avoids catastrophic performance degradation even when the Phase 1 pre-trained model is suboptimal. Furthermore, our method demonstrates significant training efficiency. For instance, the FM model trained for 300K batches with OAT-FM refinement matches the performance of the stronger OT-CFM baseline trained for the whole 400K batches (FID 3.63 vs. 3.64), and the I-CFM model at 300K with refinement (FID 3.44) outperforms the fully converged OT-CFM baseline (FID 3.44 vs. 3.64).
Table 9: Sensitivity analysis of OAT-FM to Phase 1 model quality on CIFAR-10. We evaluate the efficacy of OAT-FM fine-tuning (for only 1K batches) applied to FM, I-CFM, and OT-CFM checkpoints trained for varying durations (100K to 400K batches).
100K Batches 200K Batches 300K Batches 400K Batches
Method $downarrow NFE\downarrow FID\downarrow$ $downarrow NFE\downarrow FID\downarrow$
FM 140 6.11 140 4.26 143 3.88 147 3.71
+ OAT-FM (1K) 135 5.60 132 3.96 134 3.63 135 3.54
I-CFM 140 5.97 140 4.13 140 3.81 149 3.67
+ OAT-FM (1K) 131 5.60 137 3.95 134 3.44 138 3.48
OT-CFM 138 6.23 133 4.40 134 3.93 132 3.64
+ OAT-FM (1K) 128 6.02 126 4.18 128 3.71 126 3.46
@@ C.4 ImageNet $\times256$
C.4.1 Basic Settings
Dataset. We extend our evaluation to class-conditional generation on the ImageNet $\times256$ benchmark (deng2009imagenet). This dataset, a standard for large-scale image generation, consists of approximately 1.28 million training images, categorized into 1,000 classes.
Training. Our generative model for ImageNet is a Scalable Interpolant Transformer (SiT) (ma2024sit), which utilizes the Diffusion Transformer (DiT) (peebles2023scalable) backbone. We downloaded the checkpoint from https://github.com/willisma/SiT. The model operates in the latent space of a pre-trained variational autoencoder (VAE) (kingma2013auto). The network architecture is the XL/2 version of SiT. The interpolant framework allows for flexible choices of path-type and model prediction targets, with linear path and velocity prediction being the default configuration. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and no weight decay. The target network is updated via an EMA strategy of the online model’s weights with a decay of 0.9999. For the OAT-FM objective, we set the balancing hyperparameter $\alpha$ to 0.80. When applying Algorithm 1, to prevent cross-class interference within mini-batches, our sampling strategy assigns a unique class to each GPU for every training iteration. This ensures that each local batch consists solely of images from a single class. To incorporate Classifier-Free Guidance (CFG), we follow the training protocol of SiT-XL and randomly drop class labels with a probability of 0.1. This procedure partitions each mini-batch into two subsets: a conditional (labeled) group and an unconditional (unlabeled) group. Our method computes the OAT plan and performs the corresponding sample pairing independently within each subset. Finally, the paired samples from both groups are combined to compute the training loss. The pre-trained SiT-XL is trained by OAT-FM with additional 48K iterations or 5 epochs. Training performance metrics (including memory and training time) are detailed in Table 10.
Table 10: Training efficiency and resource consumption of SiT-XL with and without OAT-FM across different GPU configurations.
Model GPU Batchsize Peak Allocated / Wall-clock Training Time
Reserved Memory Speed 5 epochs
SiT-XL A6000 $\times$ 8 $128\times 8$ 35.57 GB / 44.61 GB 276 samples/s $\sim$6.1 hours
++ OAT-FM A6000 $\times$ 8 $128\times 8$ 35.61 GB / 44.62 GB 225 samples/s $\sim$7.5 hours
SiT-XL A100 $\times$ 8 $128\times 8$ 35.57 GB / 44.49 GB 584 samples/s $\sim$2.9 hours
++ OAT-FM A100 $\times$ 8 $128\times 8$ 35.61 GB / 44.49 GB 440 samples/s $\sim$3.8 hours
SiT-XL A100 $\times$ 8 $256\times 8$ 55.67 GB / 72.61 GB 614 samples/s $\sim$2.7 hours
++ OAT-FM A100 $\times$ 8 $256\times 8$ 55.91 GB / 72.83 GB 452 samples/s $\sim$3.7 hours
@@ C.4.1 Basic Settings
Evaluations. We employ both ODE and SDE samplers for evaluations. For ODE-based sampling, we utilize the adaptive step-size Dopri5 solver (dormand1980family), configured with an absolute tolerance of $1\times 10^{-6}$ and a relative tolerance of $1\times 10^{-3}$. For SDE-based sampling, we use a fixed-step Euler-Maruyama solver with 250 steps. In both settings, we leverage classifier-free guidance to improve sample quality.
C.4.2 More Experiment Results
The results achieved under different CFG scales are shown in Table 11. In addition, we generate images of different classes by the original SiT-XL and that refined by OAT-FM, respectively. For each example, we start from the same noise point to generate images by the two models. Following existing methods (ma2024sit; peebles2023scalable), we set the CFG scale to be 4.0 for good visual effects. Some typical results are shown in Figures 6 and 8, demonstrating that the refinement achieved by OAT-FM indeed leads to better image quality.